OpenStudy (anonymous):

simplify. Express in simplest radical form. 7√6•5√3

OpenStudy (jdoe0001):

\(\bf 7\sqrt{6}\cdot 5\sqrt{3}\implies 7\cdot 5\sqrt{6\cdot 3}\)

OpenStudy (anonymous):

Thank you. I knew it was simple. I have terrible text book.

OpenStudy (jdoe0001):

yw